OpenStudy (anonymous):

Evaluate the expression for the given values of the variables. - (4h - j) / j+h ; h = -2 and j=3

OpenStudy (anonymous):

Please help I don't understand how to solve?

OpenStudy (anonymous):

Just plug in the values -(4((-2))-3)/3-2 -(-8-3)/3-2 -(-11)/3-2 11/3-2 11/1 FINAL ANSWER IS 11

OpenStudy (anonymous):

That's it! wow thanks alot got so confused on when to distribute the negative sign!

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

I have one more similar question I do not get on my homework! |-4b - 8| + -1-b2| + 2b3 ; b= -2

OpenStudy (anonymous):

this is almost the same thing, but you have absolute value, " | |" which mean anything inside will be always positive, for example, |-3| would turn in to 3 because of the absolute value. Try the problem now, and if you still cant get it, i will help you solve it.

OpenStudy (anonymous):

| - 4(-2) -8| + | -1 - (4)2| + 2(4)3 |-8 -8| + | -1 - 16| + 128 |-16| + |-17| + 128 -33+ 128 = 95?

OpenStudy (anonymous):

oh what +33 + 128 = 161?

OpenStudy (anonymous):

it would be 24 instead of 128 so the answer would 33+24= 57

OpenStudy (anonymous):

i think you did this wrong a bit, so i will show you the way i did it, and see if you can figure mistake

OpenStudy (anonymous):

umm i dont know how to type the powers but it was b2 imeant b to the second power so it would be 16 right and 2b3 as 4 to the theird power than multiplied by 2 would make it 128 right?

OpenStudy (anonymous):

then you did it right, good job :)

OpenStudy (anonymous):

thanks a ton finally understandin this stuff!!!!!!!!!!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

i have on extra credit problem you don't have to solve, nut if you have the time i do not get it at all! Solve the equation, Check for extraneous solutions. (students are not expected to know the material) 9| 9-8x| = 2x+3